Two functions f and g are defined on the set of real numbers, R, by
f : x = x/2-x
g : x = px + q, where p and q are constants.
(i) Find g^−1
(ii) Given that g^−1(4) = 2 and g ◦ f(4) = −8, find the values of p and q
1
Expert's answer
2020-07-06T18:06:20-0400

i)g-1=inverting g(x)

 g(x)=px+q

y-q=px

g-1="(x-q)\/p"


ii) Taking g ◦ f(4) = −8 to get the first equation;

g(4/2-4)=-8

 g(-2)=-8

p(-2)+q=-8

-2p+q=-8…….(i)


Taking g^-1 to get the second equation;

Y=(x-q)/p where y=2;

2p=4-q

2p+q=4………(ii)


Taking the two simultaneous equations and solving by elimination method;

-2p+q=-8

 2p+q= 4

-4p=-12


Hence p=3 and q=-2